construct the cumulative frequency distribution for the given data

+ Share nsert Design Layout References Mailings Review Matt T 5216453318questionid 1&flushedsfalse&cld 5329723&centerwineyes Matt Theis 1/31/19 12:15 PM pring 2019 Math 1080-50(0ntro to Stat) D2L Linked Homework: Practice Exam 1(b) of Math 1080-50 Spring 2019 Save Hw Score: 0.51%,0.19 of 38 pts 3 of 28 (2 complete) Score: 0 of 1 pt Question Help 2.1.21 Age (years) of Best Actress when award was won Lesse an 30 Cumulative Frequency Construct the cumulative frequency distribution for the given data 35 30-39 Less than 50 Less than 60 Less than 70 Less than 80 70-79 ic Enter your andwer in the edit Selds a and then click Check rticde firit wihurAll parts showing MacBook Pro G Search or type URL 8 9 0 delete
Answer #1

Solution:

Given that,

Class Frequency Cumulative Frequency
20 - 29 27 27
30-39 35 27 + 35 =63
40 - 49 11 63 +11 = 74
50 - 59 2 74 + 2 = 76
60 - 69 6 76 + 6 = 82
70 - 79 1 82 +1 = 83
80 - 89 1 83 + 1 = 84
Total 84